McCann Catching, Inc. has 2.00 million shares of stock outstanding. The stock currently sells for $12.62 per share. The firm’s debt is publicly traded and was recently quoted at 90.00% of face value. It has a total face value of $10.00 million, and it is currently priced to yield 10.00%. The risk free rate is 4.00% and the market risk premium is 8.00%. You’ve estimated that the firm has a beta of 1.17. The corporate tax rate is 38.00%. The firm is considering a $43.70 million expansion of their production facility. The project has the same risk as the firm overall and will earn $12.00 million per year for 7.00 years. What is the WACC for McCann Catching, Inc.?
Answer #1

Shares outstanding = 2,000,000

Share price = $12.62

Market value of equity = Number of shares*Share price = 2,000,000*12.62

Market value of equity = $25240000

The required rate of return R(e) is calculated by CAPM model

R(e) = r(f) + Beta*(R(m) - r(f))

R(m) is the market return =8%

r(f) is the risk-free rate = 4%

R(e) = 0.04+1.17*(0.08-0.04)

R(e) = 0.0868 = 8.68%

Market value of debt = Face value*Price

Market value of debt = 10,000,000*0.90 = 9,000,000

Total value = Equity + Debt = 9,000,000+25240000 = 34240000

Weight of debt = 9000000/34240000 = 0.2629

Weight of equity= 25240000/34240000 = 0.7371

WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=
0.2629*0.11*(1-0.38)+0.7371*0.0868 = 0.01792978+0.06398028
WACC = 0.0819 = 8.19%